OpenStudy (anonymous):

Given the values n = 20, p = 0.7, and m = 14, compute the probability of m successes in n independent trials of a binary experiment.

OpenStudy (anonymous):

@Hero @saifoo.khan @tcarroll010 @Agent_Sniffles help?

OpenStudy (anonymous):

P ( m = 14 ) = 20C14 * .7^14 * . 3^6

OpenStudy (anonymous):

so it would be 20C14 * 0.7^14 * 0.3^6 = 38760 *0.7^14 * 0.3^6 = 0.191638983